OpenStudy (anonymous):

help..I don't understand how to do this algebra problem. (5, 9); x+2y=3

OpenStudy (saifoo.khan):

Plug the values in that equation and check if u are the answer 3 or not.

OpenStudy (anonymous):

Hi Saifoo..hang on a sec..

OpenStudy (saifoo.khan):

Hey Island girl. sure.

OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. Express your answer in the form of y=mx+b...sorry, i should have put what it was calling for...

OpenStudy (anonymous):

the equation of the line y=_______

OpenStudy (saifoo.khan):

Alright, for that u have find the slope for the first line. coz when it says parallel, it means that u are dealing with lines of the same slope. so the slope of the first lines = x+2y=3 y = -1/2 x + 3 slope = -1/2 now use this to find the equation of the line with the points given (5, 9) y - y1 = m (x - x1) y - 9 = -1/2 (x - 5)

OpenStudy (saifoo.khan):

try solving..

OpenStudy (anonymous):

i am trying..when i tried doing problem i got as far as the slope, after i got lost and how they determined to come up with the final answer...

OpenStudy (saifoo.khan):

now use this to find the equation of the line with the points given (5, 9) y - y1 = m (x - x1) y - 9 = -1/2 (x - 5) y - 9 = -0.5x + 2.5 y = -0.5x + 2.5 + 9 y = -0.5x + 10.5

OpenStudy (saifoo.khan):

Do u want the answer in fraction?

OpenStudy (anonymous):

yes,it says to use integers or fractions for any numbers in the expression...but also want to know exactly how to come about the final answer..

OpenStudy (anonymous):

i confused..

OpenStudy (saifoo.khan):

i showed u step by step.

OpenStudy (saifoo.khan):

where are u confused?

OpenStudy (anonymous):

that last part (5, 9) y - y1 = m (x - x1) y - 9 = -1/2 (x - 5) y - 9 = -0.5x + 2.5 y = -0.5x + 2.5 + 9 y = -0.5x + 10.5 i don't know how to answer this to come up with the equation of the line y=__

OpenStudy (saifoo.khan):

there is a formula to find the equation, y - y1 = m (x - x1) in that we just have to write the cordinates and the slope.

OpenStudy (anonymous):

oh ok..can you show me how to do that>? I have 3 more word problems like this, and with this help you give me here, i can figure my last 3 with this example

OpenStudy (saifoo.khan):

i showed u already, i inserted the cordinates on their places. and the slope on it's place.

OpenStudy (saifoo.khan):

try reading the whole thing again..

OpenStudy (anonymous):

i will..thanks again Saifoo.

OpenStudy (saifoo.khan):

No problem. i can help u again if u are stuck.

OpenStudy (anonymous):

i am trying to figure in on paper with this.. thanks

OpenStudy (saifoo.khan):

Cool.

OpenStudy (anonymous):

y=-1/2x+23/2

OpenStudy (saifoo.khan):

correct. =D

OpenStudy (anonymous):

yay, thank you..wish i could give you another medal :(

OpenStudy (saifoo.khan):

you welcome.

OpenStudy (saifoo.khan):

=)

OpenStudy (saifoo.khan):

Later may be.

OpenStudy (anonymous):

;)

OpenStudy (saifoo.khan):

[;

OpenStudy (anonymous):

Hey Saifoo..thank you soo much for you help and for the example of this math problem, i really getting the hang of it..really wish i could give you another medal....anyways, aloha.